OpenStudy (anonymous):

What is the solution set for |2x - 3| = 11?

OpenStudy (anonymous):

2x - 3 = 11 2x = 11+3 x = 11 + 3/2 x =14/2 x = 7

OpenStudy (anonymous):

Thank You :3